# A Self-Supervised Learning Approach with Differentiable Optimization for UAV Trajectory Planning

### v1

@@ Abstract
While Unmanned Aerial Vehicles (UAVs) have gained significant traction across various fields, path planning in 3D environments remains a critical challenge, particularly under size, weight, and power (SWAP) constraints. Traditional modular planning systems often introduce latency and suboptimal performance due to limited information sharing and local minima issues. End-to-end learning approaches streamline the pipeline by mapping sensory observations directly to actions but require large-scale datasets, face significant sim-to-real gaps, or lack dynamical feasibility. In this paper, we propose a self-supervised UAV trajectory planning pipeline that integrates a learning-based depth perception with differentiable trajectory optimization. A 3D cost map guides UAV behavior without expert demonstrations or human labels. Additionally, we incorporate a neural network-based time allocation strategy to improve the efficiency and optimality. The system thus combines robust learning-based perception with reliable physics-based optimization for improved generalizability and interpretability. Both simulation and real-world experiments validate our approach across various environments, demonstrating its effectiveness and robustness. Our method achieves a 31.33% improvement in position tracking error and 49.37% reduction in control effort compared to the state-of-the-art.
@@ I Introduction
Over the past decade, interest in Unmanned Aerial Vehicles (UAVs) has grown rapidly across a wide range of fields, including applications such as 3D mapping [1], exploration [2], physical interaction [3, 4, 5] and package delivery [6, 7]. One of the critical UAV tasks is efficient path planning. In the environment without a pre-established map, the UAVs must re-plan quickly and efficiently to avoid collisions and perform safe navigation even under size, weight, and power (SWAP) constraints. Path planning for UAVs poses unique challenges comparing to the planning for ground mobile robot. First, UAVs operate in three-dimensional (3D) space, significantly increasing the search area and requiring consideration of obstacles at varying altitudes. Second, UAVs typically move at higher speeds, imposing stricter safety requirements. Third, planned paths must be dynamically feasible to account for 3D maneuvers, avoiding unplanned collisions resulting from discrepancies between planned and executed paths.
@@
Recent research has explored hybrid approaches that combine learning-based and model-based methods by integrating differentiable traditional components with learning modules. For example, [14] introduced a self-supervised method combining a learning-based depth perception module with trajectory optimization (TO), using a pre-built traversability cost map for local planning. However, this approach is limited to 2D ground robots due to its reliance on a 2D cost map projection. Moreover, it employs a closed-form solution for cubic spline in TO without gradient backpropagation as claimed Bi-level optimization (BLO), reducing its generalizability to more complex or constrained scenarios. while [15] extends [14] by adding kinematic constraints, it is still limited to 2D scenario and only deals with equality constraints.
To address these limitation, we propose a self-supervised pipeline for UAV path planning that combines a learning-based depth perception with differentiable, metric-based TO, forming a bi-level optimization (BLO). Our approach includes a 3D cost map to account for UAVs’ 3D operations, providing traversability costs to guide behavior without requiring expert demonstrations or human labels. We also develop a differentiable minimum snap TO module to ensure dynamically feasible trajectories with both equality and inequality constraints and enable gradient backpropagation on iterative optimization for end-to-end training. Additionally, a time allocation network predicts segment durations, enhancing efficiency and optimality. During deployment, the policy predicts traversability conditions directly from first-person view (FPV) depth observations and plans paths accordingly. The end-to-end design optimizes observation features for planning objectives, combining the robustness of learning-based perception with the reliability, generalizability, and interpretability of physics-based methods.
@@
  • We introduce a 3D cost map that provides traversability costs to guide UAV behavior for self-supervision without the need for expert demonstrations or human labels.
  • We develop a differentiable minimum snap TO module for dynamically feasible trajectories while allowing gradient backpropagation during end-to-end training. A time allocation network enhances efficiency and optimality.
@@ II-A UAV Path Planning
Many classic planning frameworks have demonstrated effective for autonomous UAV navigation in challenging environments. For example, the modular approach [10] demonstrated in the DARPA Subterranean Challenge introduces latency and lacks robustness to noise and real-world errors. Gradient-based planning approaches optimize trajectories with manually designed safety constraints but can get stuck in local minima [16, 17, 18] . Similarly, corridor-based methods rely on low-dimensional search algorithms, often neglecting higher-order kinematics and producing dynamically infeasible trajectories [18, 19]. End-to-end learning methods have become emerging by eliminating explicit mapping and reducing latency. For instance, [11] trained a deep neural network (DNN) using human flight data, but it requires large datasets and struggles to generalize to diverse environments, with performance limited by the suboptimality of expert trajectories. Reinforcement learning has also been explored for UAV path planning [12], using real-world data to overcome the sim-to-real gap [13]. However, these approaches rely heavily on neural networks and lack guarantees for kinematic feasibility and trajectory optimality.
Recently, hybrid methods combining neural networks with numerical optimization are emerging. While many work still treated the two separately [20], there are some research incorporates differentiable optimization for end-to-end training [21, 22, 23]. For example, [21] embeds safe corridors and trajectory optimization into neural networks that can efficiently generate collision-free and dynamically feasible trajectories for UAVs; however, it relies on supervised learning with results limited to simulations. [23] uses an end-to-end strategy integrating perception and trajectory optimization by predicting the offsets and scores of a set of motion primitives for local optimization. However, the sampling-based method and predicted end-derivative cannot guarantee optimality. [22] integrate DNNs within the motion planning framework to predict the time allocation of a given set of waypoints. However, it formulates path searching as a separate module and rely on supervised learning for training.
@@ II-B Differentiable Optimization
Differentiable optimization refers to optimization algorithms where the output is differentiable with respect to the input parameters, enabling the integration of traditional model-based optimization with learning-based approaches. This paradigm combines the strengths of both methods—infusing interpretable domain knowledge into deep learning pipelines and reducing the need for tedious parameter tuning in traditional optimization. It has become a powerful paradigm in robotics for perception [24], planning [25], control [26], and physics-based simulation [27]. A pioneering work, OptNet [28], embeds optimization as a differentiable layer within neural networks, enabling gradient backpropagation for argmin problems. This approach has later been extended to model predictive control (MPC) using convex quadratic approximations for non-convex MPC problems [29], and to actor-critic MPC by integrating RL [30]. SDPRLayers further expanded this to semidefinite program relaxations for non-convex polynomial problems [31]. Recently, BPQP reformulated the backward pass as a quadratic programming (QP) problem [32], improving computational efficiency. Tools like Theseus [33] and PyPose [34] have also emerged, offering differentiable optimization capabilities for robotics. Our approach builds on our previous work [15, 35, 34], formulating the problem as a BLO where gradient backpropagation from the upper to lower level is critical, and employs a differentiable QP solver to incorporate both equality and inequality physical constraints.
@@ III-B Planning with Hybrid Learning and Optimization
The proposed hybrid learning and optimization process for UAV planning is shown in Figure 2 and consists of three key modules. First, a convolutional neural network (CNN) front-end takes the depth input and encode into observation embedding. This embedding, combined with the target, is fed into the second planning network to predict a n-key-point path $\bm{\xi}\in\mathbb{R}^{n\times 3}$. In the back end, a Time Allocation Net takes into the key-point path and generates the timestamp $\mathcal{T}\in\mathbb{R}^{n}$ for each trajectory segment. Next, a differentiable minimum snap trajectory optimizer (MSTO) takes into both the key-point path, generated time allocation and outputs a dynamic feasible trajectory $\bm{\tau}^{*}$. Then, a well-designed traversability costs $\mathcal{U}$ based on a 3D cost map is evaluated and backpropagated through both the differentiable MSTO and the networks to update the network parameters of both the front-end network $\bm{\theta}_{f}$ and the Time Allocation Net $\bm{\theta}_{T}$. This process results in a nested bi-level optimization (BLO) problem, where the networks and differentiable MSTO can be jointly optimized.
@@ III-D1 UAV model
where $m$ is the UAV mass, $\mathbf{r}=[x,y,z]^{\top}$ is the UAV position vector in world frame $\mathcal{W}$, $g$ is the gravity vector, $F$ is the body thrust force magnitude, $\mathbf{J}$ is the moment of inertia matrix along the body principle axes. $\bm{\omega}$ is the robot angular velocity, $\mathbf{m}=[m_{x},m_{y},m_{z}]$ is the body moment. We use 3-2-1 Euler angles $\bm{\Omega}=[\phi,\theta,\psi]$ to define the orientation. The system state is defined as: $\mathbf{x}=[\mathbf{r},\bm{\Omega},\dot{\mathbf{r}},\bm{\omega}]^{\top}$, and control input $\mathbf{u}=[F;\mathbf{m}]^{\top}$
@@ III-D2 Trajectory Generation
It has been proved that the quadrotor dynamics is differential flat with the flat outputs as $\bm{\sigma}=[\mathbf{r}^{\top},\psi]^{\top}\in\mathbb{R}^{4}~{}$[37]. In other words, the states and the inputs can be written as algebraic functions of the UAV position and heading and their derivatives. Therefore, the trajectory planning can be simplified as planning for $\bm{\sigma}$, which also ensures the dynamic feasibility. In fact, the inputs $m_{x}$ and $m_{y}$ appear as functions of the $4^{\mathrm{th}}$ derivatives of the positions (snap), $m_{z}$ appears in the $2^{\mathrm{nd}}$ derivative of the yaw angle, the TO problem is formulated as minimizing the total snap and control effort along a time horizon $t_{n}$ given key-point time stamp as $\mathcal{T}=[t_{1},...,t_{n}]^{\top}\in\mathbb{R}^{n}$:
@@
where $\mu_{r}$ and $\mu_{\psi}$ are constant weight factors to make the integral nondimensional. $\bm{\sigma}(t)$ are piecewise polynomials of order $m$ over $n$ (Here $m=7$) time intervals for each dimension of the flat output $\mathbf{r}(t)$ and $\psi(t)$ parameterized by time $t$. $\bm{\sigma}_{i}(t)=[\mathbf{r}(t)_{i}^{\top},\psi(t)_{i}]^{\top}$ denotes the polynomials of the $i^{\mathrm{th}}$ segment between time interval $t_{i}$ and $t_{i+1}$. $(\cdot)^{[k]}$ represents the quantity up to $k^{\mathrm{th}}$ order derivative, e.g $\mathbf{r}^{[k]}(t)=[\mathbf{r}(t)^{\top},\mathbf{r}^{\prime}(t)^{\top},...,\mathbf{r}^{(k)}(t)^{\top}]^{\top}$. $(\cdot)_{0},(\cdot)_{T}$ are initial and final boundary conditions. Specifically, (3b) constrains the waypoint and its higher-order derivatives at the start and end points, the intermediate waypoints is constrained by (3c), while (3d) ensures the continuity of the position and its higher-order derivatives at the intermediate waypoints. Notice that the hard constraints [37] can be relaxed to inequality constraints to incorporating flight corridor, actuator limit, etc.
In fact, we can represent $\bm{\sigma}_{i}=\mathbf{c}_{i}^{\top}\bm{\Omega}(t)$, where $\bm{\Omega}(t)=[1,t,...,t^{m}]^{\top}$. With the coefficients of all the piecewise polynomials as the decision variables $\mathbf{c}=[\mathbf{c}_{1}^{\top},...,\mathbf{c}_{n}^{\top}]^{\top}$, Problem (3) can be formulated as a QP problem
@@
s.t. $\mathbf{A}\mathbf{c}=\mathbf{b}$
$\mathbf{G}\mathbf{c}\leq\mathbf{h}$
Notice that the key point path $\bm{\xi}$ is predicted by the front-end network planner $f(\bm{\theta}_{f})$. Problem (4) is the inner/lower level optimization or Problem (1). With optimized $\mathbf{c}^{*}$, we then evaluate $\bm{\sigma}(t)$ at control frequency resulting the planned trajectory $\bm{\tau}^{*}$.
@@ III-D3 Time Allocation
In Problem (3), the arrival times at different key points are critical but cannot be predetermined, especially in dynamic environments. A common approach in previous work is to solve a separate iterative optimization problem using gradient descent with backtracking line search [37]. However, this will lead to a tri-level optimization in our setup, significantly degrading real-time performance. To address this, we develop a Time Allocation Net $g(\bm{\theta_{T}})$ that predicts the time allocation for each segment of a key-point trajectory. It takes the key-point path $\bm{\xi}$ as input and output the allocated times $\mathcal{T}$. We use the traditional gradient-descent-based method to compute optimal time allocations for training. Then, during inference, our network efficiently predicts the time allocations in real time. In practice, the Time Allocation Net utilizes an MLP followed by a softmax layer, predicting the time allocation percentage for each trajectory segment relative to the total time.

### v2

@@ Abstract
While Unmanned Aerial Vehicles (UAVs) have gained significant traction across various fields, path planning in 3D environments remains a critical challenge, particularly under size, weight, and power (SWAP) constraints. Traditional modular planning systems often introduce latency and suboptimal performance due to limited information sharing and local minima issues. End-to-end learning approaches streamline the pipeline by mapping sensory observations directly to actions but require large-scale datasets, face significant sim-to-real gaps, or lack dynamical feasibility. In this paper, we propose a self-supervised UAV trajectory planning pipeline that integrates a learning-based depth perception with differentiable trajectory optimization. A 3D cost map guides UAV behavior without expert demonstrations or human labels. Additionally, we incorporate a neural network-based time allocation strategy to improve the efficiency and optimality. The system thus combines robust learning-based perception with reliable physics-based optimization for improved generalizability and interpretability. Both simulation and real-world experiments validate our approach across various environments, demonstrating its effectiveness and robustness. Our method achieves a 30.90% reduction in control effort while maintaining competitive tracking performance compared with state-of-the-art.
@@ I Introduction
Over the past decade, interest in Unmanned Aerial Vehicles (UAVs) has grown rapidly across a wide range of fields, including applications such as 3D mapping [35], exploration [13], physical interaction [11, 9] and package delivery [8]. One of the critical UAV tasks is efficient path planning. In the environment without a pre-established map, the UAVs must re-plan quickly and efficiently to avoid collisions and perform safe navigation even under size, weight, and power (SWAP) constraints. Path planning for UAVs poses unique challenges comparing to the planning for ground mobile robot. First, UAVs operate in three-dimensional (3D) space, significantly increasing the search area and requiring consideration of obstacles at varying altitudes. Second, UAVs typically move at higher speeds, imposing stricter safety requirements. Third, planned paths must be dynamically feasible to account for 3D maneuvers, avoiding unplanned collisions resulting from discrepancies between planned and executed paths.
@@
Recent research has explored hybrid approaches that combine learning-based and model-based methods by integrating differentiable traditional components with learning modules [30]. For example, [32] introduced a self-supervised method combining a learning-based depth perception module with trajectory optimization (TO), using a pre-built traversability cost map for local planning. However, this approach is limited to ground robots due to its reliance on a 2D cost map projection. Moreover, it employs a closed-form solution for cubic spline in TO without gradient backpropagation as claimed Bi-level optimization (BLO), reducing its generalizability to more complex or constrained scenarios. While [18] extends [32] with differentiable model predictive control (MPC) to enforce kinematic constraints, it guarantees only kinematic (not dynamic) feasibility due to its bicycle-model state choice. It is also limited to a 2D setting and handles only equality constraints. [10] embeds safe corridors and differentiable TO in the network for generating UAV dynamically feasible trajectories. However, it requires an offline motion-primitive library and still relies on supervised training.
To address these limitation, we propose a self-supervised pipeline for UAV path planning that combines a learning-based depth perception with differentiable, metric-based TO, forming a bi-level optimization (BLO). Our approach includes a 3D cost map to account for UAVs’ 3D operations, providing collision costs to guide behavior without requiring expert demonstrations or human labels. We also develop a differentiable minimum snap TO module to ensure dynamically feasible trajectories with both equality and inequality constraints and enable gradient backpropagation on iterative optimization for end-to-end training. Additionally, a time allocation network predicts segment durations, enhancing efficiency and optimality. During deployment, the policy predicts collision conditions directly from first-person view (FPV) depth observations and plans paths accordingly. The end-to-end design optimizes observation features for planning objectives, combining the robustness of learning-based perception with the reliability, generalizability, and interpretability of physics-based methods.
@@
  • We train the planner using geometry-derived collision signals from a 3D cost map for self-supervision without the need for expert demonstrations or human labels.
  • We develop a differentiable minimum snap TO module for dynamically feasible trajectories while allowing end-to-end training. A time allocation network is designed to enhance efficiency and optimality.
@@ II-A UAV Path Planning
Many classic planning frameworks have demonstrated effective for autonomous UAV navigation in challenging environments. For example, the modular approach [6] demonstrated in the DARPA Subterranean Challenge introduces latency and lacks robustness to noise and real-world errors. Gradient-based planning approaches optimize trajectories with manually designed safety constraints but can get stuck in local minima [37, 38, 7]. [36] is a UAV motion planning framework that leverages Euclidean Distance Fields (EDF) for gradient-based B-spline optimization. Similarly, corridor-based methods rely on low-dimensional search algorithms, often neglecting higher-order kinematics and producing dynamically infeasible trajectories [7, 28]. End-to-end learning methods have become emerging by eliminating explicit mapping and reducing latency. For instance, [19] trained a deep neural network (DNN) using human flight data, but it requires large datasets and struggles to generalize to diverse environments, with performance limited by the suboptimality of expert trajectories. Reinforcement learning has also been explored for UAV path planning [17], using real-world data to overcome the sim-to-real gap [16]. However, these approaches rely on neural networks and lack guarantees for kinematic feasibility and trajectory optimality.
Recently, hybrid methods combining neural networks with numerical optimization are emerging. While many work still treated the two separately [14], there are some research incorporates differentiable optimization for end-to-end training [10, 31, 20]. For example, [20] integrates perception and trajectory optimization by predicting the offsets and scores of a set of motion primitives for local optimization. However, the sampling-based method and predicted end-derivative cannot guarantee optimality. [31] integrate DNNs within the motion planning framework to predict the time allocation of a given set of waypoints. However, it formulates path searching as a separate module and rely on supervised learning for training. In contrast, our method avoids handcrafted primitives and uses self-supervision to jointly optimize perception and planning for 3D UAV scenario, reducing labeling and integration overhead.
@@ II-B Differentiable Optimization
Differentiable optimization refers to optimization algorithms where the output is differentiable with respect to the input parameters, enabling the integration of traditional model-based optimization with learning-based approaches. This paradigm combines the strengths of both methods—infusing interpretable domain knowledge into deep learning pipelines and reducing the need for tedious parameter tuning in traditional optimization. It has become a powerful paradigm in robotics for perception [27], planning [32], control [15], and physics-based simulation [34]. A pioneering work, OptNet [3], embeds optimization as a differentiable layer within neural networks, enabling gradient backpropagation for argmin problems. This approach has later been extended to model predictive control (MPC) using convex quadratic approximations for non-convex MPC problems [2], and to actor-critic MPC by integrating RL [26]. Recently, BPQP reformulated the backward pass as a quadratic programming (QP) problem [22], improving computational efficiency. Tools like Theseus [23] and PyPose [29] have also emerged, offering differentiable optimization capabilities for robotics. Our approach formulates the problem as a BLO where gradient backpropagation from the upper to lower level is critical, and employs a differentiable QP solver to incorporate both equality and inequality physical constraints.
@@ III-B Planning with Hybrid Learning and Optimization
The proposed hybrid learning and optimization process for UAV planning is shown in Figure 2. First, front-end network $f$: a convolutional neural network (CNN) front-end takes the depth input and encode into observation embedding. This embedding, combined with the target, is fed into the second planning network to predict a n-key-point path $\bm{\xi}\in\mathbb{R}^{n\times 3}$. In the back end, a Time Allocation Net $g$ takes into the key-point path and generates the timestamp $\mathcal{T}\in\mathbb{R}^{n}$ for each trajectory segment. Next, a differentiable minimum snap trajectory optimizer (MSTO) takes into both the key-point path, generated time allocation and outputs a dynamic feasible trajectory $\bm{\tau}^{*}$. Then, a well-designed collision costs $\mathcal{U}$ based on a 3D cost map is evaluated and backpropagated through both the differentiable MSTO and the networks to update the network parameters of both the front-end network $\bm{\theta}_{f}$ and the Time Allocation Net $\bm{\theta}_{T}$. This process results in a nested bi-level optimization (BLO) problem, where the networks and differentiable MSTO can be jointly optimized.
@@ III-D1 UAV model
where $m$ is the UAV mass, $\mathbf{r}=[x,y,z]^{\top}$ is the UAV position vector in world frame $\mathcal{W}$, $g$ is the gravity, $F$ is the body thrust, $\mathbf{J}$ is the moment of inertia along the body principle axes. $\bm{\omega}$ is the robot angular velocity, $\mathbf{m}=[m_{x},m_{y},m_{z}]$ is the body moment. We use 3-2-1 Euler angles $\bm{\Omega}=[\phi,\theta,\psi]$ to define the orientation. The system state is defined as: $\mathbf{x}=[\mathbf{r},\bm{\Omega},\dot{\mathbf{r}},\bm{\omega}]^{\top}$, and control input $\mathbf{u}=[F;\mathbf{m}]^{\top}$.
@@ III-D2 Trajectory Generation
It has been proved that the quadrotor dynamics is differential flat with the flat outputs as $\bm{\sigma}=[\mathbf{r}^{\top},\psi]^{\top}\in\mathbb{R}^{4}$~[21]. In other words, the states and the inputs can be written as algebraic functions of the UAV position and heading and their derivatives. Therefore, the trajectory planning can be simplified as planning for $\bm{\sigma}$ rather than the full dynamic states, which also ensures the dynamic feasibility. Computation efficiency can thus be improved compared to the traditional nonlinear MPC. In fact, the inputs $m_{x}$ and $m_{y}$ appear as functions of the $4^{\mathrm{th}}$ derivatives of the positions (snap), $m_{z}$ appears in the $2^{\mathrm{nd}}$ derivative of the yaw angle, the TO problem is formulated as minimizing the total snap and control effort along a time horizon $t_{n}$ given key-point time stamp as $\mathcal{T}=[t_{1},...,t_{n}]^{\top}\in\mathbb{R}^{n}$:
@@
where $\mu_{r}$ and $\mu_{\psi}$ are constant weight factors to make the integral nondimensional. $\bm{\sigma}(t)$ are piecewise polynomials of order $m$ over $n$ (Here $m=7$) time intervals for each dimension of the flat output $\mathbf{r}(t)$ and $\psi(t)$ parameterized by time $t$. $\bm{\sigma}_{i}(t)=[\mathbf{r}_{i}(t)^{\top},\psi_{i}(t)]^{\top}$ denotes the polynomials of the $i^{\mathrm{th}}$ segment between time interval $t_{i}$ and $t_{i+1}$. $(\cdot)^{[k]}$ represents the quantity up to $k^{\mathrm{th}}$ order derivative, e.g $\mathbf{r}^{[k]}(t)=[\mathbf{r}(t)^{\top},\mathbf{r}^{\prime}(t)^{\top},...,\mathbf{r}^{(k)}(t)^{\top}]^{\top}$. $(\cdot)_{0},(\cdot)_{T}$ are initial and final boundary conditions. Specifically, (3b) constrains the waypoint and its higher-order derivatives at the start and end points, the intermediate waypoints is constrained by (3c), while (3d) ensures the continuity of the position and its higher-order derivatives at the intermediate waypoints. Notice that the hard constraints can be relaxed to inequality constraints to incorporating flight corridor, actuator limit, etc.
In fact, we can represent $\bm{\sigma}_{i}=\mathbf{c}_{i}^{\top}\bm{\Omega}(t)$, where $\bm{\Omega}(t)=[1,t,...,t^{m}]^{\top}$. With the coefficients of all the piecewise polynomials as the decision variables $\mathbf{c}=[\mathbf{c}_{1}^{\top},...,\mathbf{c}_{n}^{\top}]^{\top}$, Problem (3) can thus be formulated as a QP problem:
@@
s.t. $\mathbf{A}\mathbf{c}=\mathbf{b},\quad\mathbf{G}\mathbf{c}\leq\mathbf{h}$
where $\mathbf{A}$, $\mathbf{b}$ encapsulate the equality constraints, and $\mathbf{G}$, $\mathbf{h}$ describe the inequality constraints. Notice that the key point path $\bm{\xi}$ is predicted by the front-end network planner $f(\bm{\theta}_{f})$. Problem (4) is the inner/lower level optimization of Problem (1). With optimized $\mathbf{c}^{*}$, we then evaluate $\bm{\sigma}(t)$ at control frequency to obtain the planned trajectory $\bm{\tau}^{*}$.
@@ III-D3 Time Allocation
In Problem (3), the arrival times at different key points are critical but cannot be predetermined, especially in dynamic environments. A common approach in previous work is to solve a separate iterative optimization problem using gradient descent with backtracking line search [21]. However, this will lead to a tri-level optimization in our setup, significantly degrading real-time performance. To address this, we develop a Time Allocation Net (TAN) $g(\bm{\theta_{T}})$ that predicts the time allocation for each segment of a key-point trajectory. It takes the key-point path $\bm{\xi}$ as input and output the allocated times $\mathcal{T}$. We use the traditional gradient-descent-based method as in  [21] to compute optimal time allocations for training. Then, during inference, our network efficiently predicts the time allocations in real time. In practice, TAN utilizes an MLP followed by a softmax layer, predicting the time allocation percentage for each trajectory segment relative to the total time. Notice that TAN was jointly trained with the front-end network with the supervised reference generated by the optimization method [21], rather than expert demonstration or human annotations.